OpenStudy (anonymous):

Parametrize a disk of radius a which lies on the yz plane and is oriented clockwise when viewed from the positive x-axis.

OpenStudy (anonymous):

I'm not too sure about what you mean by 'oriented clockwise'. I'm going to assume you mean that the position vector rotates clockwise when we view down along the positive x-axis. If the disk lies in the y-z plane, with y horizontal and z vertical, then the position vector will be, \[(y,z)=({a}\cos(-{\theta}),{a}\sin(-{\theta}))=({a}\cos({\theta}),-{a}\sin({\theta}))\]Setting

OpenStudy (anonymous):

\[\theta \rightarrow -{\theta}\]will allow the position vector to rotate clockwise as you increase theta. I hope this is what you're looking for.

OpenStudy (anonymous):

Yes, thank you.